OpenStudy (anonymous):

SOMEONE HELP PLEASE!! Suppose a Baseball player had 247 hits in a season. In the given probability distribution, the random variable X represents the number of hits the player obtained in the game. (a). Compute and interpret the mean of the random variable X. x , P(x) 0 , 0.1427 1 , 0.4898 2 , 0.2612 3 , 0.0615 4 , 0.0222 5 , 0.0226

OpenStudy (kropot72):

The mean value of X is calculated from \[\sum_{}^{}xP(x)=(0\times 0.1427)+(1\times 0.4898)+(2\times 0.2612)+.........\]